# Topological transition in filamentous cyanobacteria: from motion to structure

### v2

@@ I Introduction
Figure 1: Dynamics of gliding motility. (a) Close-up micrograph of Oscillatoria lutea filaments, composed of chains of individual cells. The scale bar is $30~{}\mu$m. (b) A time lapse of three micrographs demonstrates filament motion, by which each filament’s body glides along the track laid down by its head; the time lapse between each frame is 51 s, the scale bars in each frame are $100~{}\mu$m. The rightmost image highlights this path-tracking dynamics by overlaying the three images using shadings of different colors. (c) Schematic depiction of the model filaments, as discretized into chains of connected beads. Beads are placed at a fixed distance $v_{i}\delta t$ apart. Upon advancing time by an increment $\delta t$, the last bead of a filament is removed and placed as its new head. Non-reciprocal interactions occur when a filament’s head (filament $i$) is within a distance $\sigma$ (pink disk; not drawn to scale) from any bead of another filament, $j$. The head experiences a nematic aligning interaction $\mathcal{F}(\theta_{i})$ while filament $j$ remains unaffected by this interaction.
@@
These gliding cyanobacteria are becoming recognized as the archetype of a distinct class of active filaments [72, 73, 74, 75, 76]. Unlike active polymers, their motion cannot be represented as resulting from active forces tangential to monomer-like elements [49, 77]. Gliding cyanobacteria exert negligible forces on each other, as opposed to the stresses that they apply on their substrate, which represents a very different physical case than active nematics based on microtubule-kinesin or actin-myosin [78]. The physical properties that set gliding cyanobacteria apart from other active filament systems, and which help define their own class of active matter, active spaghetti, include: (i) a large aspect ratio, (ii) gliding motility induced by polar forces, (iii) path-tracking dynamics, where the body of the filament follows the track laid down by its head (also known as metameric locomotion [52]), and (iv) non-reciprocal aligning interactions.
@@
Here, we explore the dynamics, structure, and phase behavior of a system of active filaments, modeled after the behavior of cyanobacteria. We investigate the large-scale collective effects of cyanobacteria colonies, while still retaining information about the individual constituents’ dynamics and their interactions. Combining numerical simulations with theoretical arguments, we uncover a rich dynamics for this model of gliding filaments as their Péclet number $\mathrm{Pe}$, area coverage $\Phi$, and filament length $L$ are varied. As it develops towards a statistically steady state, this model shows transient behaviors that closely resemble corresponding experiments with filamentous cyanobacteria. We also find that the filaments’ effective translational diffusivity grows with the Péclet number, $\mathrm{Pe}$. The nonequilibrium phase diagram of the system exhibits a transition from isotropic to reticulate patterns as Péclet number and area coverage, or density, increase. Finally, inspired by polymer physics, we compute the parallel and perpendicular pair correlation function for the filaments. The features of these correlation functions carries information about the structure of the reticulate network and faithfully predicts its characteristic length scale.
Taken together, our results identify the curvature fluctuations and aligning interactions as the mechanism underpinning the nonequilibrium transition to the reticulate pattern in filamentous cyanobacteria.
@@ II.1 Model for gliding cyanobacteria
Figure 2: Time evolution of systems of filaments forming into reticulate structures. (a)-(e) micrographs of experiments performed with Oscilatoria lutea, where $v_{0}=3\,\mu$m/s and $\tau=8\,$min [72]. (f)-(k) simulations of a colony initialized with randomly oriented filaments, using a filament length $L=1$, a Péclet number $\mathrm{Pe}=3.5$ and an area coverage $\Phi=1$. The field of view shown and time between snapshots match the experimental measurements of O. lutea (i.e. $t=2\,\mathrm{h}/\tau=15$ corresponds to 2 hours). The scale bars in the experimental images are 101010\,10mm and scale bars in simulations 101010\,10mm$/(v_{0}\tau)$. In both experiments and simulations the initial pattern forms with smaller features that coarsen slightly to achieve a stable length scale over a period of a few hours ($t\approx 50$).
@@
The last term in Eq. (4) describes the interactions, of strength $J$, between nearby filaments, where $\mathcal{F}(\theta_{i})=N_{ij}^{-1}\frac{\partial}{\partial\theta_{i}}\sum_{i\sim j}\cos(\theta_{i}-\theta_{j}^{\bar{\alpha}})$ is a Lebwohl–Lasher interaction [84, 85] inducing nematic alignment between the filaments, $N_{ij}$ is the number of filaments within an interaction distance $\sigma$ of the $i$-th filament’s head, $\sum_{i\sim j}$ denotes the sum over these filaments, and the orientation angle $\theta_{j}^{\bar{\alpha}}$ of the bead closest to the $i$-th head $\mathbf{x}_{i}$. The interaction has a short, finite range $\sigma$, which is consistent with the filaments’ effective diameter; when the distance between the head of filament $i$ and any bead of filament $j$ is less than $\sigma$, the orientation of the head of filament $i$ undergoes a deflection due to the Lebwohl–Lasher interaction, however, the bead of filament $j$ remains unperturbed. This makes the interactions fundamentally non-reciprocal, and it is predicated on the experimental observation that when two trichomes meet, only the incident filament’s head participates in the alignment process, while the other filament is not deflected [72]. Non-reciprocal interactions explicitly break detailed balance [86, 87] and are common in living active matter [88]. A schematic depiction of the non-reciprocal nematic interactions used in our model is shown in Fig. 1(c).
@@
Given the equation of motion (3) and (4), a Péclet number may be defined as $\mathrm{Pe}=v_{0}/(l\sqrt{D_{\omega}\tau})$ with a length scale $l$. This dimensionless number may be understood as a ratio of how quickly curvature $\kappa$ is transported within the system compared with the time it takes to randomize a filament’s curvature.
Figure 5: Emergence of reticulate patterns. The mosaic plot shows the steady-state behavior of the system at different area coverage $\Phi=0.25,0.5,0.75,1.0,1.25$ and Péclet numbers $\mathrm{Pe}=0.875,1.75,3.5,7.0,14.0$. The individual panels show simulation snapshots at steady state in a $30\times 30$ domain with periodic boundary conditions. Filament length is kept constant at $L=1$. The panels’ frames indicate the system’s average nematic order $\langle S\rangle$, calculated as an average of $1\times 1$ sized blocks. With increasing $\mathrm{Pe}$, the system transitions from a disordered state into one showing reticulate patterns of nematic bundles. The inset at $(\mathrm{Pe},\Phi)=(3.5,1.0)$ shows an example of the spatially resolved nematic order parameter $S_{\text{loc}}$; regions of higher density correlate with higher local nematic order parameter. Bundles show a high degree of ordering, whereas dilute regions in between are less ordered.
@@ III.1 Dynamics and transient
We first investigate the dynamical aspects of the collective self-organization of cyanobacteria filaments. At sufficiently large number density $\rho$ and Péclet number $\mathrm{Pe}$ trichomes self-organize into reticulate patterns [61, 72] as shown in Fig. 2. We study here the process by which this reticulate pattern emerges from an isotropic initial condition, and compare model predictions of the transient dynamics with experimental data. In simulations, we prepare a system of filaments at area coverage $\Phi=1$ with a homogeneous distribution of filaments positions and isotropic distribution of their orientations, see Fig. 2. For the corresponding experiments, images of the colony dynamics were cropped and contrast-enhanced to show the central 17$\times$17 mm2 of the well plate, and are given in Fig. 2.
In both experiments and simulations, the aligning interactions lead to the formation of short, dense bundles of filaments. Over time, these bundles connect and self-organize into a network topology. During an initial transient period, the typical separation length between bundles, which provides a characteristic length scale for the patterns, slowly increases. After a few hours in the experiments, or simulation times of $t\sim 50$, the pattern then converges into a steady-state configuration of reticulate bundles of high local density, with less dense regions in between. The comparison of the transient dynamics between experiments and simulations reveals that our model quantitatively reproduces the emergence of the reticulate pattern in time; in fact, for $\tau=8$ min, as measured in Ref. [72] a simulated time $t=75$ corresponds to 10 hours, which allows for a direct, one-to-one comparison to the experimental time scale, as in Fig. 2.
@@
Figure 3(b) also shows that being ordered into bundles has a small but noticeable effect on their mobility, since switching off the interactions among filaments, i.e. setting $g=0$ (which removes the possibility to form bundles) causes a small increase in their effective long-time translational diffusion $D_{\text{eff}}$. In other words, the presence of the reticulate pattern alters the motility of the filaments in a subtle way but without disrupting it; that implies that the filaments do not remain confined within any particular bundle, or loop.
@@
While the filaments’ gliding motility is of polar nature, the nematic nature of their interactions results in the nematic symmetry within the bundles, with filaments moving along them in either direction in equal parts, see Fig. 4. The approximate equal distribution of orientations within bundles was also measured experimentally in Ref. [72]. While the resulting patterns may look reminiscent of vortex lattices observed in other active systems, the filaments are not constrained to vortices and explore the paths formed by the network of bundles.
@@ III.2 Structure
Figure 8: Emergence of reticulate patterns. The mosaic plot shows the steady-state collective behavior of the system for different filament lengths $L$ and Péclet numbers $\mathrm{Pe}$. The individual panels show simulation snapshots at steady state in a $30\times 30$ domain with periodic boundary conditions. The density of simulations is set to ensure an equal area coverage, such that $\Phi=1$. The panels’ frames indicate the system’s average nematic order $\langle S\rangle$, calculated as an average of $1\times 1$ sized blocks. With increasing $\mathrm{Pe}$, the system transitions from a disordered state into one showing reticulate patterns of nematic bundles. The length scale of the emerging patterns is not affected by filament length.
@@
The second dimensionless parameter governing the filaments’ dynamics is the interaction strength $g$. We explore how varying $g$ affects the structure formation in Fig. 7. In the limit of $g\to 0$, the system remains isotropic. Upon increasing $g$, the reticulate pattern emerges ($g\approx 2$). This is the parameter range that most closely resembles the experimental observations of young O. lutea colonies. For larger values of $g$, $g\approx 4$, the interactions dominate over the curvature fluctuations, and we observe aster-like structures (see e.g.  [91, 92], though in a different physical context), and strong coils for $g\gtrsim 8$ (see e.g. [93]).
@@
We can obtain more precise information about the effects of the filament length $L$ in our model cyanobacteria system by considering the limiting behavior as $L$ approaches zero. Figure 9 shows how the average nematic order parameter $\langle S\rangle$ depends on the Péclet number for different $L$. Generally, upon increasing $\mathrm{Pe}$ the system transitions from an isotropic state, characterized by a vanishingly small value of $\langle S\rangle$, to a reticulate pattern, characterized by $\langle S\rangle\simeq 0.7$. The transition point can be identified with the inflection point of the curve at $\mathrm{Pe}=\mathrm{Pe_{c}}\simeq 1.8$, which does not significantly change with filament length $L$. The spatial extent of the filaments effectively introduces memory into the system. So, while the chance of interacting with another filament remains the same, for a constant $\Phi$, the result of an aligning interaction will be retained within the system for a time $L/v_{0}$. This fact enables longer filaments to retain some degree of local order, even if the system is isotropic at larger scales. The nature of the observed phase transition thus changes with increasing filament length. Short filaments experience a sharper transition from disordered to ordered states while longer filaments experience a much smoother transition around this point. In the limit $L\to 0$, the system for $\Phi=1$ appears to approach a second-order or continuous transition at an apparent critical point $\mathrm{Pe_{c}}\simeq 1.8$. The characterization of the critical scaling in the proximity of $\mathrm{Pe_{c}}$ is deferred to future work.
@@
Figure 10: Emergence of structure. Filament head-head pair correlation functions (PCFs) parallel to the direction of motion, $g_{\parallel}(r)$ (dashed lines), and orthogonal to it, $g_{\perp}(r)$ (solid lines), see Eqs. (9)-(10), quantify the structure of the reticulate pattern. (a) The time evolution (color axis) of these functions is shown for a system with area coverage $\Phi=1$, filament length $L=1$ and Péclet number $\mathrm{Pe}=3.5$. As time advances and the reticulate pattern forms, $g_{\parallel}(r)$ and $g_{\perp}(r)$ converge to a steady state. For a system with a distribution of curvatures $\delta\kappa$, the minimum of $g_{\perp}(r)$ coincides with $\mathrm{Pe}=(v_{0}\tau\delta\kappa)^{-1}$, indicated by an arrow on the $x$-axis. (b) Steady-state PCFs are shown for different $\mathrm{Pe}$, keeping filament length $L=1$ and $\Phi=1$. The minima in $g_{\perp}(r)$ coincide with the length scales predicted from the distributions of curvatures, indicated by color-coded arrows. The shoulder in $g_{\parallel}(r)$ moves to larger distances with increasing $\mathrm{Pe}$ (except for $\mathrm{Pe}=9.9$, where finite size effects are visible). (c) Effect of filament length $L$ (color bar) on the PCFs at $\mathrm{Pe}=3.5$. The number density $\rho$ is scaled here, to fix $\Phi=1$. Independent of $L$, the minimum in $g_{\perp}(r)$ coincides with the system’s $Pe$, as denoted by an arrow. The magnitude of the extrema in $g_{\parallel}$ and $g_{\perp}$ decrease with increasing filament length, indicating less well-defined patterns. (d) Temporal evolution of the typical length scale $\ell^{*}$ associated to the reticulate pattern, obtained from the minimum of $g_{\perp}(r)$ in panel (a). At long times, $\ell^{*}\to 3.5$, which is the value of $\mathrm{Pe}$. (e) Dependence of $\ell^{*}$ on $\mathrm{Pe}$ showing that the minimum of $g_{\perp}(r)$ matches the Péclet number and is a good measure of the reticulate’s typical length scale; the dashed line corresponds to the identity.
A powerful method to obtain structural information in soft-matter systems is the pair correlation function (PCF) $g(r)$ [94]. This is defined as the probability of finding another particle within a volume $dV$ at distance $r$ from a reference particle. In polymer physics, it is possible to compute the PCF for any pair of monomers within the same chain or between different chains [95]. In our system, a simple choice is to consider the head-head PCF, where only the head bead of each filament is considered. Furthermore, given the long persistence length of the filament it is convenient to consider $g(r)$ along the directions parallel and normal to the direction of motion of the filament’s head. We thus define the head-head parallel PCF

### v3

@@ I Introduction
Figure 1: Dynamics of gliding motility. (a) Close-up micrograph of Oscillatoria lutea filaments, composed of chains of individual cells. The scale bar is $30~{}\mu$m. (b) A time lapse of three micrographs demonstrates filament motion, by which each filament’s body glides along the track laid down by its head; the interval between frames is 51 s; the scale bars in each frame are $100~{}\mu$m. The rightmost image highlights this path-tracking dynamics by overlaying the three images using shadings of different colors. (c) Schematic depiction of the model filaments, as discretized into chains of connected beads. Beads are placed at a fixed distance $v_{i}\delta t$ apart. Upon advancing time by an increment $\delta t$, the last bead of a filament is removed and placed as its new head. Non-reciprocal interactions occur when a filament’s head (filament $i$) is within a distance $\sigma$ (pink disk; not drawn to scale) from any bead of another filament, $j$. The head experiences a nematic aligning interaction $\mathcal{F}(\theta_{i})$ while filament $j$ remains unaffected by this interaction.
@@
These gliding cyanobacteria are becoming recognized as the archetype of a distinct class of active filaments [72, 73, 74, 75, 76]. Unlike active polymers, their motion cannot be represented as resulting from active forces tangential to monomer-like elements [49, 77]. Gliding cyanobacteria exert negligible forces on each other, as opposed to the stresses that they apply on their substrate; thus, physical interactions are different than active nematics based on microtubule-kinesin or actin-myosin [78]. The physical properties that set gliding cyanobacteria apart from other active filament systems, and which help define their own class of active matter, active spaghetti, include: (i) a large aspect ratio, (ii) gliding motility induced by polar forces, (iii) path-tracking dynamics, where the body of the filament follows the track laid down by its head (also known as metameric locomotion [52]), and (iv) non-reciprocal aligning interactions.
@@
Here, we explore the dynamics, structure, and phase behavior of a system of active filaments, modeled after the behavior of cyanobacteria. We investigate the large-scale collective effects of cyanobacteria colonies, while still retaining information about the individual constituents’ dynamics and their interactions. Combining numerical simulations with theoretical arguments, we uncover a rich dynamics for this model of gliding filaments as their Péclet number $\mathrm{Pe}$, area coverage $\Phi$, and filament length $L$ are varied. As it develops towards a statistically steady state, this model shows transient behaviors that closely resemble corresponding experiments with filamentous cyanobacteria. We also find that the filaments’ effective translational diffusivity grows with the Péclet number, $\mathrm{Pe}$. The nonequilibrium phase diagram of the system exhibits a transition from isotropic to reticulate patterns as the Péclet number and area coverage, or density, increase. Finally, inspired by polymer physics, we compute the parallel and perpendicular pair correlation functions for the filaments. The features of these correlation functions carries information about the structure of the reticulate network and faithfully predicts its characteristic length scale.
Taken together, our results identify the curvature fluctuations and aligning interactions as the mechanisms underpinning the nonequilibrium transition to the reticulate pattern in filamentous cyanobacteria.
@@ II.1 Model for gliding cyanobacteria
Figure 2: Time evolution of systems of filaments forming into reticulate structures. (a)-(e) Micrographs of experiments performed with Oscilatoria lutea, where $v_{0}=3\,\mu$m/s and $\tau=8\,$min [72]. (f)-(k) Simulations of a colony initialized with randomly oriented filaments, using a filament length $L=1$, a Péclet number $\mathrm{Pe}=3.5$ and an area coverage $\Phi=1$. The field of view shown and time between snapshots match the experimental measurements of O. lutea (i.e., simulated time $t/\tau=15$ corresponds to 2 h). The scale bars in the experimental images are 101010\,10mm and scale bars in simulations 101010\,10mm$/(v_{0}\tau)$. In both experiments and simulations the initial pattern forms with smaller features that coarsen slightly to achieve a stable length scale over a period of a few hours ($t\approx 50$).
@@
The last term in Eq. (4) describes the interactions, of strength $J$, between nearby filaments, where $\mathcal{F}(\theta_{i})=N_{ij}^{-1}\frac{\partial}{\partial\theta_{i}}\sum_{i\sim j}\cos(\theta_{i}-\theta_{j}^{\bar{\alpha}})$ is a Lebwohl–Lasher interaction [84, 85] inducing nematic alignment between the filaments, $N_{ij}$ is the number of filaments within an interaction distance $\sigma$ of the $i$-th filament’s head, $\sum_{i\sim j}$ denotes the sum over these filaments, and the orientation angle $\theta_{j}^{\bar{\alpha}}$ refers to the bead closest to the $i$-th head $\mathbf{x}_{i}$. The interaction has a short, finite range $\sigma$, which is consistent with the filaments’ effective diameter; when the distance between the head of filament $i$ and any bead of filament $j$ is less than $\sigma$, the orientation of the head of filament $i$ undergoes a deflection due to the Lebwohl–Lasher interaction, however, the bead of filament $j$ remains unperturbed. This makes the interactions fundamentally non-reciprocal, and it is predicated on the experimental observation that when two trichomes meet, only the incident filament’s head participates in the alignment process, while the other filament is not deflected [72]. Non-reciprocal interactions explicitly break detailed balance [86, 87] and are common in living active matter [88]. A schematic depiction of the non-reciprocal nematic interactions used in our model is shown in Fig. 1(c).
@@
Given the equations of motion (3) and (4), a Péclet number may be defined as $\mathrm{Pe}=v_{0}/(l\sqrt{D_{\omega}\tau})$ with a length scale $l$. This dimensionless number may be understood as a ratio of how quickly curvature $\kappa$ is transported within the system, compared with the time it takes to randomize a filament’s curvature.
Figure 5: Emergence of reticulate patterns. The mosaic plot shows the steady-state behavior of the system at different area coverage $\Phi=0.25,0.5,0.75,1.0,1.25$ and Péclet numbers $\mathrm{Pe}=0.875,1.75,3.5,7.0,14.0$. The individual panels show simulation snapshots at steady state in a $30\times 30$ domain with periodic boundary conditions. Filament length is kept constant at $L=1$. The panels’ frames indicate the system’s average nematic order $\langle S\rangle$, calculated as an average of $1\times 1$ sized blocks. With increasing $\mathrm{Pe}$, the system transitions from a disordered state into one showing reticulate patterns of nematic bundles. The inset at $(\mathrm{Pe},\Phi)=(3.5,1.0)$ shows an example of the spatially resolved nematic order parameter $S_{\text{loc}}$; regions of higher density correlate with higher local nematic order. Bundles show a high degree of ordering, whereas dilute regions in between are less ordered.
@@ III.1 Dynamics and transient
We first investigate the dynamical aspects of the collective self-organization of cyanobacteria filaments. At sufficiently large number density $\rho$ and Péclet number $\mathrm{Pe}$ trichomes self-organize into reticulate patterns [61, 72] as shown in Fig. 2. We study here the process by which this reticulate pattern emerges from an isotropic initial condition, and compare model predictions of the transient dynamics with experimental data. In simulations, we prepare a system of filaments at area coverage $\Phi=1$ with a homogeneous distribution of filament positions and isotropic distribution of their orientations, see Fig. 2. For the corresponding experiments, images of the colony dynamics were cropped and contrast-enhanced to show the central 17$\times$17 mm2 of the well plate, and are given in Fig. 2.
In both experiments and simulations, the aligning interactions lead to the formation of short, dense bundles of filaments. Over time, these bundles connect and self-organize into a network topology. During an initial transient period, the typical separation length between bundles, which provides a characteristic length scale for the patterns, slowly increases. After a few hours in the experiments, or simulation times of $t\gtrsim 50$, the pattern then converges into a steady-state configuration of reticulate bundles of high local density, with less dense regions in between. The comparison of the transient dynamics between experiments and simulations reveals that our model quantitatively reproduces the emergence of the reticulate pattern in time; in fact, for $\tau=8$ min, as measured in Ref. [72] a simulated time $t=75$ corresponds to 10 hours, which allows for a direct, one-to-one comparison to the experimental time scale, as in Fig. 2.
@@
Figure 3(b) also shows that the ordering of filaments into bundles has a small but noticeable effect on their mobility, since switching off the interactions among filaments, i.e. setting $g=0$ (which removes the possibility to form bundles) causes a small increase in their effective long-time translational diffusion $D_{\text{eff}}$. In other words, the presence of the reticulate pattern alters the motility of the filaments in a subtle way but without disrupting it; that implies that the filaments do not remain confined within any particular bundle, or loop.
@@
While the filaments’ gliding motility is of polar nature, the nematic nature of their interactions results in the nematic symmetry within the bundles, with filaments moving along them in either direction in equal parts, see Fig. 4. The approximate equal distribution of orientations within bundles was also measured experimentally in Ref. [72]. While the resulting patterns may look reminiscent of vortex lattices observed in other active systems [91], the filaments are not constrained to vortices and explore the paths formed by the network of bundles.
@@ III.2 Structure
Figure 8: Emergence of reticulate patterns. The mosaic plot shows the steady-state collective behavior of the system for different filament lengths $L$ and Péclet numbers $\mathrm{Pe}$. The individual panels show simulation snapshots at steady state in a $30\times 30$ domain with periodic boundary conditions. The filament density is set to ensure an equal area coverage, such that $\Phi=1$. The panels’ frames indicate the system’s average nematic order $\langle S\rangle$, calculated as an average of $1\times 1$ sized blocks. With increasing $\mathrm{Pe}$, the system transitions from a disordered state into one showing reticulate patterns of nematic bundles. The length scale of the emerging patterns is not affected by filament length.
@@
The second dimensionless parameter governing the filaments’ dynamics is the interaction strength $g$. We explore how varying $g$ affects the structure formation in Fig. 7. In the limit of $g\to 0$, the system remains isotropic. Upon increasing $g$, the reticulate pattern emerges ($g\approx 2$). This is the parameter range that most closely resembles the experimental observations of young O. lutea colonies. For larger values, $g\approx 4$, the interactions dominate over the curvature fluctuations, and we observe aster-like structures (see e.g.  [92, 93], though in a different physical context), and strong coils for $g\gtrsim 8$ (see e.g. [94]).
@@
We can obtain more precise information about the effects of the filament length $L$ in our model cyanobacteria system by considering the limiting behavior as $L$ approaches zero. Figure 9 shows how the average nematic order parameter $\langle S\rangle$ depends on the Péclet number for different $L$. Generally, upon increasing $\mathrm{Pe}$ the system transitions from an isotropic state, characterized by a vanishingly small value of $\langle S\rangle$, to a reticulate pattern, characterized by $\langle S\rangle\simeq 0.7$. The transition point can be identified with the inflection point of the curve at $\mathrm{Pe}=\mathrm{Pe_{c}}\simeq 1.8$, which does not significantly change with filament length $L$. The spatial extent of the filaments effectively introduces memory into the system. So, while the chance of interacting with another filament remains the same, for a constant $\Phi$, the result of an aligning interaction will be retained within the system for a time $L/v_{0}$. This fact enables longer filaments to retain some degree of local order, even if the system is isotropic at larger scales. The nature of the observed phase transition thus changes with increasing filament length. Short filaments experience a sharper transition from disordered to ordered states while longer filaments experience a much smoother transition around this point. In the limit $L\to 0$, the system for $\Phi=1$ appears to approach a continuous transition at an apparent critical point $\mathrm{Pe_{c}}\simeq 1.8$. The characterization of the critical scaling in the proximity of $\mathrm{Pe_{c}}$ is deferred to future work.
@@
Figure 10: Emergence of structure. Filament head-head pair correlation functions (PCFs) parallel to the direction of motion, $g_{\parallel}(r)$ (dashed lines), and orthogonal to it, $g_{\perp}(r)$ (solid lines), see Eqs. (9)-(10), quantify the structure of the reticulate pattern. (a) The time evolution (color axis) of these functions is shown for a system with area coverage $\Phi=1$, filament length $L=1$ and Péclet number $\mathrm{Pe}=3.5$. As time advances and the reticulate pattern forms, $g_{\parallel}(r)$ and $g_{\perp}(r)$ converge to a steady state. For a system with a distribution of curvatures $\delta\kappa$, the minimum of $g_{\perp}(r)$ coincides with $\mathrm{Pe}=(v_{0}\tau\delta\kappa)^{-1}$, indicated by an arrow on the $x$-axis. (b) Steady-state PCFs are shown for different $\mathrm{Pe}$, keeping filament length $L=1$ and $\Phi=1$. The minima in $g_{\perp}(r)$ coincide with the length scales predicted from the distributions of curvatures, indicated by color-coded arrows. The shoulder in $g_{\parallel}(r)$ moves to larger distances with increasing $\mathrm{Pe}$ (except for $\mathrm{Pe}=9.9$, where finite size effects are visible). (c) Effect of filament length $L$ (color bar) on the PCFs at $\mathrm{Pe}=3.5$. The number density $\rho$ is scaled here to fix $\Phi=1$. Independently of $L$, the minimum in $g_{\perp}(r)$ coincides with the system’s Pe, as denoted by an arrow. The magnitude of the extrema in $g_{\parallel}$ and $g_{\perp}$ decrease with increasing filament length, indicating less well-defined patterns. (d) Temporal evolution of the typical length scale $\ell^{*}$ associated to the reticulate pattern, obtained from the minimum of $g_{\perp}(r)$ in panel (a). At long times, $\ell^{*}\to 3.5$, which is the value of $\mathrm{Pe}$. (e) Dependence of $\ell^{*}$ on $\mathrm{Pe}$ showing that the minimum of $g_{\perp}(r)$ matches the Péclet number and is a good measure of the reticulate’s typical length scale; the dashed line corresponds to the identity.
A powerful method to obtain structural information in soft-matter systems is the pair correlation function (PCF) $g(r)$ [95]. This is defined as the probability of finding another particle within a volume $dV$ at distance $r$ from a reference particle, relative to a homogeneous distribution. In polymer physics, it is possible to compute the PCF for any pair of monomers within the same chain or between different chains [96]. In our system, a simple choice is to consider the head-head PCF, where only the head bead of each filament is considered. Furthermore, given the long persistence length of the filament it is convenient to consider $g(r)$ along the directions parallel and normal to the direction of motion of the filament’s head. We thus define the head-head parallel PCF